OpenStudy (tiffany_rhodes):

solve for x:

OpenStudy (tennis5518):

where the number

OpenStudy (kayders1997):

What's the question?

OpenStudy (tiffany_rhodes):

\[\sum_{n=1}^{nifty} (2x)^(11n) =44\]

OpenStudy (tiffany_rhodes):

11n is the power

OpenStudy (anonymous):

nifty?

OpenStudy (freckles):

it looks like you try to write (2x)^(11n) ?

OpenStudy (tiffany_rhodes):

yeah, sorry. I'm not used to the formatting. It's a sum from n=1 to infinity of 2x raised to the power 11n equal to 44

OpenStudy (freckles):

\[\sum_{n=1}^{\infty} 2x^{11n}=44 \text{ or } \sum_{n=1}^{\infty} (2x)^{11n}=44\] either way we will use geometric series thingy formula :p

OpenStudy (tiffany_rhodes):

Yeah, so that's what I was thinking. Divide the 2 over and you get x^(11n) = 22 and the sum of a geometric series = a1/(1-r) where r is the common ratio. I know that a geometric series only converges (and has a sum) if the common ratio r < 1.

OpenStudy (freckles):

\[\sum_{n=1}^{\infty} r^{k}=\frac{r}{1-r}\]

OpenStudy (freckles):

where r would be ?

OpenStudy (tiffany_rhodes):

x ... ?

OpenStudy (tiffany_rhodes):

That's the part where I was confused.

OpenStudy (anonymous):

are there any restrictions on the value of x?

OpenStudy (tiffany_rhodes):

yeah, x < 1

OpenStudy (tiffany_rhodes):

if x= r

OpenStudy (anonymous):

i mean in the question

OpenStudy (anonymous):

oops sorry didnt read the question correctly

OpenStudy (tiffany_rhodes):

oh, no. Not that I'm aware of. I included all of the information about the question

OpenStudy (freckles):

almost \[2 \sum_{n=1}^{\infty} x^{11n}=2 \sum_{n=1}^{\infty} (x^{11})^n\]

OpenStudy (freckles):

x^(11) should be r

OpenStudy (tiffany_rhodes):

oh okay.

OpenStudy (freckles):

also I messed up my formula earlier I put n's and k's :p should be n's or k's not both \[\sum_{n=1}^{\infty} r^{n}=\frac{r}{1-r} \\ \sum_{n=1}^{\infty} (x^{11})^n=\frac{?}{1-?}\]

OpenStudy (tiffany_rhodes):

Alright, I got x = (22/23)^(1/11) and it was correct. Thanks so much @freckles :)

OpenStudy (freckles):

np :)

OpenStudy (tennis5518):

Oh wow you are a smarty @freckles